OpenStudy (anonymous):

I need some help these discrete math questions: For each part below, find the number of passwords that fit the condition. (a) The password has 10 alphanumeric characters, but is not case-sensitive. (b) The password has 10 to 12 alphanumeric characters, and is case-sensitive. (c) The password has 10 case-sensitive alphanumeric characters, and exactly 3 of them are numerical. (d) The password has 10 case-sensitive alphanumeric characters, exactly 3 of them are numerical, and no numerical character is repeated.

OpenStudy (anonymous):

a) I found like 36^10 but I'm not sure..

OpenStudy (amistre64):

if alphanumeric defines just letters and numbers, then yes, there are 36 options for each position.

OpenStudy (amistre64):

26*2 = 52 + 10 = 62, there are 62 options for each position with some modifications i believe for the b part

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Alphanumeric

OpenStudy (anonymous):

I thought that in that way..

OpenStudy (amistre64):

62^10+62^11+62^12 is what im thinking for b such that abcdefghij differs from 0abcdefghij, 00abcdefghij

OpenStudy (anonymous):

what does 10 to 12 alphanumeric characters mean, can we choose 10 alphanumeric, 11 alphanumeric or 12 alphanumeric ?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

hmm, it makes sense now..

OpenStudy (anonymous):

for c 52^7*10^3

OpenStudy (amistre64):

for c im thinking: N,N,N,a,a,a,a,a,a,a given that N has 10 and a has 52 \[10C3~(10^3+52^7)\]but that wouldnt account for 333 being the same as 333 would it

OpenStudy (amistre64):

im over thinking that one ...

OpenStudy (anonymous):

I guess my solution is just valid for if first three passwords is number....right..

OpenStudy (amistre64):

yes, since there is not a position defined, we would have to take that value and multiply it by the number of ways the positions can change

OpenStudy (amistre64):

the last one is the same as the one before it, we just have to subtract the set of "same numbers" NNN NNn NnN nNN there appears to be 4 ways that that set can be arranged, so all we would have to do is figure out how many are in the set to begin with

OpenStudy (amistre64):

001 991 010 ... 919 ; 10 ways 100 199 002 992 020 ... 929 200 299 ... 009 999 090 ... 999 900 999 ----- 10 ways is it safe to say there are 100 ways to get duplicated numbers? or am i introducing a double count?

OpenStudy (amistre64):

3*10*10 = 300 ways .... and yes there are some double counts

OpenStudy (amistre64):

i cant think thru that one :/

OpenStudy (anonymous):

can we say for c 10^35*2^7*62^10

OpenStudy (amistre64):

for c; 10^3*52^7 is valid for one positional set up there are \[\frac{10!}{3!7!}\]distinct ways to set it up

OpenStudy (anonymous):

typo 10^3

OpenStudy (amistre64):

10.9.8 = 720 ; /6 = 120 distinct ways

OpenStudy (anonymous):

but you added them previous solution, was your that solution wrong,,

OpenStudy (amistre64):

the addition was a mistake on my part ... brain fell asleep while typing :)

OpenStudy (anonymous):

yeah, brains do that sometimes (:

OpenStudy (amistre64):

001 991 010 ... 919 ; 10 ways ... but 111,111,111 suggests that we need to adjust 100 199 by subtracting 2 3*10 -2 = 28, there are 10 ways to run thru this making: c - 280 can we see any issues with this idea for "d" ?

OpenStudy (anonymous):